OpenStudy (anonymous):

What are the uses for the ribs in terrestrial vertebrates for taking air into the lungs?

thomaster (thomaster):

intercostal muscles between the ribs make the rib cage expand, that will create an underpressure in the longs so air will flow in :D

OpenStudy (anonymous):

o-O You got a source for that, sir? cx

thomaster (thomaster):

well yea Campbell Biology ninth edition :P

thomaster (thomaster):

I can try to find it on the internet if you'd like that

OpenStudy (anonymous):

If only I knew how to cite a book. Cx

OpenStudy (anonymous):

Merci beaucoup o-o

thomaster (thomaster):

\(\huge\bf\color{turquoise}{Pas~de~problème}~\color{magenta}{:D}\)

thomaster (thomaster):

And how to cite a book: \(\sf Name(year)\it book ~\sf City, State: Publisher\) So for my book that would be: Reece, J. B., Urry, L., Cain, M., Wasserman, S., Minorsky, P., & Jackson, R. (2011). Campbell biology (9th edition; International edition.). Harlow: Pearson Education

OpenStudy (anonymous):

o-O

thomaster (thomaster):

Or you can also only say the name of the first author followed by et al (which means and others) Reece, J. B., et al. (2011). Campbell biology (9th edition; International edition.). Harlow: Pearson Education

thomaster (thomaster):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (anonymous):

Woah. o-O Je suis encore a un 69 SmartScore cela prend une eternite a augmenter. :/

thomaster (thomaster):

ikr! But answering questions is faster than asking them :P

OpenStudy (anonymous):

True lol